OpenStudy (anonymous):

Two gears are connected and are rotating simultaneously. The smaller gear has a radius of 4 inches, and the larger gear has a radius of 7 inches. two circles touching at one point. Larger circle has radius of 7 inches. Smaller circle has radius of 4 inches. Part 1: What is the angle measure, in degrees and rounded to the nearest tenth, through which the larger gear has rotated when the smaller gear has made one complete rotation? Part 2: How many rotations will the smaller gear make during one complete rotation of the larger gear? Show all work

OpenStudy (dangerousjesse):

When the small circle makes one full spin, the larger one made 4/7 of a spin. One full rotation= 360 degrees, so the large one rotated \(\frac{4}{7}\times 360\)